# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v1

@@ 6.3 ReLU
Figure 3: (a) Step-by-step construction of the normalized ReLU-based local basis function from its constituent $\operatorname{ReLU}(e_{i}-x)$ and $\operatorname{ReLU}(x-s_{i})$ terms. (b) Complete set of normalized ReLU–KAN basis functions $R_{i}(x)$ with supports $s_{i},e_{i}] [$87].
@@ 6.4 Jacobi Polynomials
    where $w_{i}^{(P)}$ and $w_{j}^{(Q)}$ are trainable weights, and $\sigma(x)$ is a squashing function.
@@
In parallel, Zhang et al. [66] [^21] https://github.com/zhang-zhuo001/Legend-KINN explored the Legendre specialization $(\alpha,\beta)=(0,0)$ within Jacobi-based KANs, confirming its favorable conditioning and rapid convergence for physics-informed PDE solvers.
@@ 6.5 Gaussian (RBF)
Pure RBF Layers with Learnable Centers. A more flexible formulation introduces learnable Gaussian centers, as proposed by Abueidda et al. [46] [^23] https://github.com/DiabAbu/DeepOKAN. This allows the basis functions to adapt spatially during training. Each activation takes the form
@@
where $g_{j}\in\mathbb{R}$ are trainable centers and $\beta>0$ controls the receptive field width. The corresponding layer transformation is
@@
where $W_{l}$ are learnable synthesis weights and $G_{l}=\{g_{j}\}$ denotes the center set. By optimizing both $W_{l}$ and $G_{l}$, the network gains geometric adaptivity—allowing basis functions to migrate toward regions of high functional complexity—while preserving the smoothness and differentiability of Gaussian kernels.
@@ 6.6 Fourier
In addition to classical Fourier-based designs, Jiang et al. [43] [^32] https://github.com/Jim137/qkan propose the quantum-inspired Kolmogorov–Arnold Network (QKAN), where each univariate map is realized by a data re-uploading variational quantum circuit (QVAF). These circuits naturally generate Fourier-like expansions with exponentially scalable frequency modes, offering parameter-efficient alternatives to grid-based Fourier KANs while retaining the structured interpretability of classical KANs.
@@ 6.7 Wavelet
Learning $s$ and $u$ per neuron lets Wav-KAN adapt receptive fields and frequency resolution, aiding generalization in, e.g., image classification [103] where feature granularity varies spatially.
Patra et al. [164] integrate Wav-KAN into a physics-informed framework for coupled nonlinear PDEs, reporting faster convergence than spline-based Efficient-KAN at comparable accuracy under data-free collocation. In this setting, the wavelet activation is chosen as the sine–Gaussian
$\psi_{\text{sin}}(t)\;=\;\pi^{-\tfrac{1}{4}}\sin(\omega_{0}t)\,e^{-t^{2}/2},$
with $\omega_{0}$ controlling the center frequency.
In bioinformatics, Pratyush et al. [183] propose CaLMPhosKAN for phosphorylation-site prediction, employing a DoG wavelet. This enhances detection of residue-level signals in disordered protein regions.
For hyperspectral imaging, Seydi et al. [147] combine CWT and DWT within Wav-KAN to extract spatial–spectral features, outperforming spline-KANs and MLPs on Indian Pines and Salinas datasets.
Summary. Wavelet-based KANs provide an expressive, multiscale, and interpretable basis that improves convergence and robustness across domains—ranging from PDEs and bioinformatics to high-dimensional imaging—by uniting the locality of wavelets with the compositional structure of KANs.
@@ 7 Accuracy Improvement
To orient the reader, we begin with a compact summary of accuracy–improving mechanisms for KANs (Table 6). The table groups methods by mechanism and lists concise technique keywords with representative references. The remainder of this section expands on each group in turn, detailing when/why they help and practical choices and caveats.
Table 6: Compact map of accuracy-improving mechanisms for KANs. Cells use concise keywords; see Section 7 for context.
Mechanism Technique keywords Ref.
Physics-consistent & loss design
PIKAN Chebyshev recursion; RBA; EVM; Bayesian [29]
KAN-PINN Strongly nonlinear PDE residuals; actuator deflection [47]
KAN-MHA Navier–Stokes residuals + BC losses; attention-guided [48]
Res-KAN Residual physics + sparse regression; variable-coefficient PDEs [49]
KKAN Self-scaled residual reweighting; adaptive residual focus [44]
AL-PKAN Augmented Lagrangian; auto weight-balancing [58]
AIVT (cKAN) Velocity–vorticity loss; turbulence; residual weighting [194]
KANtrol Fractional / integro-differential operators [59]
Adaptive sampling & grids
[…]
Free-knot KAN Trainable knot positions; cumulative softmax ordering [38]
Grid extension Increase $G$ during training; optimizer state transition [82]
[…]
Multi-resolution PIKAN Scheduled coarse/fine sampling for Chebyshev-based PIKAN [64]
Domain decomposition
FBKAN (PoU) Overlapping subdomains; smooth PoU blend; local KANs [86]
Temporal subdomains Long-horizon split; improved NTK conditioning [56]
Function decomposition
MFKAN Freeze LF surrogate; linear+nonlinear HF heads; blend $\alpha$ [162]
SPIKAN Sum of products of 1D KAN factors; tensorized structure [163]
KAN-SR Recursive simplification; divide–and–conquer symbolic extraction [204]
Hybrid / Ensemble & data
MLP–KAN MLP experts for features; KAN for functional shaping [33]
HPKM-PINN Parallel MLP & KAN branches; learnable fusion [50]
KKAN Per-dim MLP features; explicit basis expansion [44]
Sequence / attention hybrids
FlashKAT (Transformer) Group-rational KAN blocks inside Transformer [166]
[…]
GINN-KAN Growing interpretable NN + KAN; PINN coupling [165]
KAN-ODE KAN as $\dot{u}$ in Neural ODE; adjoint training [36]
AAKAN-WGAN Adaptive-activation KAN + WGAN; data augmentation [168]
Discontinuities & sharp gradients
SincKAN Sinc interpolation; kinks; boundary layers [41]
rKAN Rational bases; asymptotics; jump capture [92]
DKAN $\tanh$ gate (jump) + spline background; shock localization [52]
KINN Singularities / stress concentrations; reduced spectral bias [61]
Two-phase PINN–KAN Hybrid optimizer; preserves saturation fronts [53]
Optimization & adaptive training
[…]
Bayesian Bayesian optimization of KAN hyperparameters; imbalanced data [161]
Bayesian PINN–KAN Variational inference; KL regularization; calibrated uncertainty [105]
@@ 7.1 Physics Constraints & Loss Design
  • Zhang et al. [47]: KAN-PINN, embedding strongly nonlinear PDE residuals directly into the loss, enabling accurate prediction of IPMC actuator deflections with improvements over MLP-PINNs.
  • Yang et al. [48]: KAN-MHA, integrating Navier–Stokes residuals and boundary-condition losses into the training objective to guide convergence in airfoil flow prediction.
  • Guo et al. [49]: Res-KAN, combining residual-based physics losses with sparse regression terms to recover variable-coefficient PDEs under sparse and noisy observations.
  • Shukla et al. [29]: cPIKAN + RBA + EVM, strengthening physics-informed KANs by incorporating residual-based attention and entropy-viscosity stabilization, improving robustness in high-Reynolds-number fluid simulations.
  • Toscano et al. [44]: KKAN + ssRBA, introducing self-scaled residual-based attention that adaptively reweights PDE residual losses to preserve signal quality and enhance generalization.
  • Zhang et al. [58]: AL-PKAN, replacing fixed penalty weights with learnable Lagrange multipliers through an augmented Lagrangian loss, allowing dynamic balancing among multiple physics constraints.
  • Toscano et al. [194]: AIVT, reformulating the physics loss in velocity–vorticity form to improve reconstruction of turbulent convection fields from sparse experimental inputs.
  • Aghaei [59]: KANtrol, extending the loss to fractional and integro-differential physics operators via Gaussian quadrature and operational matrices, enabling accurate handling of optimal control PDEs.
  • Chen et al. [208]: WCN, introducing a weak-form physics loss from the Fokker–Planck equation with adaptive collocation sampling, enabling efficient recovery of drift and diffusion terms in SDEs from aggregate data.
  • Gong et al. [209]: Physics-inspired KAN, embedding Navier–Stokes coupling into a multi-field K-means clustering step that constrains sensor placement and guides data-fidelity loss in indoor flow reconstruction.
  • Guo et al. [217]: PKAN, embedding prior physics into initialization and mapping spline activations to symbolic operators, enabling compact analytic equation discovery.
  • Sen et al. [212]: KAN-ETS, enforcing Ehrenfest constraints [213] in the loss to preserve quantum dynamical laws in time-series modeling.
  • Wang et al. [214]: SPEL-KAN, encoding Hamiltonian dynamics and Lie group symmetries [215] with physics-driven sparsity constraints on system matrices.
  • Lou et al. [227]: DAE-KAN, a physics-informed KAN that directly solves high-index differential–algebraic equations without index reduction, using coupled KAN branches to preserve algebraic constraints and eliminate drift-off errors.
@@ 7.2 Adaptive Sampling & Grids
Grid–adaptive PIKANs. Rigas et al. [82] extend this idea by adapting not only the spline grid but also the collocation points, so resolution follows the PDE residual:
@@ 7.3 Domain decomposition
Figure 17: Schematic of FBKAN domain decomposition and PoU blending. Overlapping subdomains $\Omega_{j}$ host lightweight local KANs $K_{j}$, smoothly combined via PoU weights $\omega_{j}(x)$. This enables parallel training and per-subdomain knot/hyperparameter choices [86].
@@ 7.4 Function decomposition
$K_{H}(x)=\underbrace{(1-\alpha)\,K_{\ell}\big(x,K_{L}(x)\big)}_{\text{linear LF}\to\text{HF trend}}+\underbrace{\alpha\,K_{nl}\big(x,K_{L}(x)\big)}_{\text{nonlinear HF correction}}\!,$
where $\alpha\in[0,1]$ is a trainable parameter, $K_{\ell}$ is the linear block, $K_{nl}$ the nonlinear block, and $K_{L}$ the frozen LF surrogate.
The loss function balances three objectives:
@@
Here, $\lambda_{\alpha}>0$ and $n\in\mathbb{N}$ control how strongly the model prefers the linear block, while $w>0$ applies weight-decay on the nonlinear block parameters $\Phi_{nl}$.
@@
In a related line of work, Jacob et al. [163] propose a separable PIKAN (“SPIKAN”), where instead of one multivariate KAN, the solution is expressed as a sum of products of 1D KAN factors (one per coordinate), which are then combined to approximate the global field.

### v2

@@ 6.3 ReLU
Figure 3: (a) Step-by-step construction of the normalized ReLU-based local basis function from its constituent $\operatorname{ReLU}(e_{i}-x)$ and $\operatorname{ReLU}(x-s_{i})$ terms. (b) Complete set of normalized ReLU–KAN basis functions $R_{i}(x)$ with supports $s_{i},e_{i}] [$123].
@@ 6.4 Jacobi and General Polynomials
  • Holomorphic Monomial Polynomials: Simple powers $(1,z,\dots,z^{P})$, where z=x+iyz=x+i\,y denotes the complex input, as used in PIHKAN [255].
@@
    where $w_{i}^{(P)}$ and $w_{j}^{(Q)}$ are trainable weights, and $\sigma(x)$ is an activation function.
@@
In parallel, Zhang et al. [74] [^23] https://github.com/zhang-zhuo001/Legend-KINN examined the Legendre specialization $(\alpha,\beta)=(0,0)$ within Jacobi-based KANs, demonstrating favorable conditioning and rapid convergence for physics-informed PDE solvers. A complementary direction is offered by the Adaptive PolyKAN [280], which treats the polynomial degree as a trainable parameter, enabling the model to adjust its functional complexity dynamically over the course of training.
@@ 6.5 Gaussian (RBF)
Pure RBF Layers with Learnable Centers. A more flexible RBF formulation appears in the DeepOKAN architecture of Abueidda et al. [54] [^25] https://github.com/DiabAbu/DeepOKAN, where Gaussian bases are used in both branch and trunk networks. Each activation takes the form
@@
with trainable centers $g_{j}$ and width $\beta>0$. The layer update is
@@
allowing the RBFs to migrate toward regions of high variation. In contrast to some later operator-learning studies, Abueidda et al. report that RBF–KANs resolve high-frequency sinusoidal fields with markedly higher accuracy than MLPs, confirming the strong approximation capability of Gaussian bases in mechanics problems.
A complementary conclusion arises in the recent two-step DeepOKAN for bubble dynamics [143]: while RBFs remain effective in the branch network, spline bases outperform them in the trunk for extremely high-frequency, high-amplitude oscillations. Together, these works show that RBFs are highly competitive for moderate-to-high frequency regimes, but basis choice in KAN operator networks remains problem-dependent.
@@ 6.6 Fourier
In addition to classical Fourier-based designs, Jiang et al. [51] [^34] https://github.com/Jim137/qkan propose the quantum-inspired Kolmogorov–Arnold Network (QKAN), where each univariate map is realized by a data re-uploading variational quantum circuit. These circuits naturally generate Fourier-like expansions with exponentially scalable frequency modes, offering parameter-efficient alternatives to grid-based Fourier KANs while retaining the structured interpretability of classical KANs.
@@ 6.7 Wavelet
Learning $s$ and $u$ per neuron allows Wav-KAN to adapt receptive fields and frequency resolution, aiding generalization in settings with spatially varying feature scales, such as image classification [139].
Patra et al. [191] apply wavelet activations—using a sine–Gaussian basis—to physics-informed training, tuning frequency response to obtain faster data-free convergence on coupled nonlinear PDEs. Meshir et al. [270] complement this with an NTK analysis, showing that the Morlet mother-wavelet frequency (or hidden-layer width) directly controls NTK eigenvalue decay, providing explicit spectral-bias mitigation and strong gains in physics-informed settings (Poisson/Heat/Wave/Helmholtz).
Applications further highlight the versatility of wavelet bases: Pratyush et al. [210] employ a DoG wavelet in CaLMPhosKAN for enhanced phosphorylation-site detection, while Seydi et al. [174] combine CWT and DWT for hyperspectral imaging, outperforming spline-KANs and MLPs.
Summary. Wavelet-based KANs provide a flexible multiscale basis whose localized frequency control improves convergence and robustness across PDEs, bioinformatics, and high-dimensional imaging.
@@ 7 Accuracy Improvement
To orient the reader, we begin with a compact summary of accuracy–improving mechanisms for KANs. The table groups methods by mechanism and lists concise technique keywords with representative references. The remainder of this section expands on each group in turn, detailing when/why they help and practical choices and caveats.
@@ 7.1 Physics Constraints & Loss Design
  • Structural & Conservation Laws. Enforce hard invariants such as Hamiltonian, quantum, or algebraic constraints. Includes SPEL-KAN [241], KAN-ETS [239], and DAE-KAN [254].
  • Weak-Form & Variational Constraints. Use integral/variational formulations instead of pointwise PDE residuals. Includes WCN [235] and KANtrol [67].
  • Alternative Problem Formulations. Modify the optimization objective or physical variables. Includes AL-PKAN [66], AIVT [221], cPIKAN+RBA+EVM [37], KKAN+ssRBA [52], KAN-MHA [56], and holomorphic–potential PIHKAN [255].
  • Physics-Guided Discovery & Design. Physics-based losses used for inference, discovery, or design tasks. Includes PKAN [244] and physics-guided sensor design [236].
  • General PDE–Residual Models. Standard PINN-style residuals without special structural constraints. Includes KAN-PINN [55], Res-KAN [57], Anant-KAN [256], and PI-KAN [269].
@@ 7.2 Adaptive Sampling & Grids
Grid–adaptive PIKANs. Rigas et al. [118, 259] extend this idea by adapting not only the spline grid but also the collocation points, so resolution follows the PDE residual:
@@ 7.3 Domain decomposition
Figure 17: Schematic of FBKAN domain decomposition and PoU blending. Overlapping subdomains $\Omega_{j}$ host lightweight local KANs $K_{j}$, smoothly combined via PoU weights $\omega_{j}(x)$. This enables parallel training and per-subdomain knot/hyperparameter choices (adapted from [122]).
@@ 7.4 Function decomposition
$K_{H}(x)=(1-\alpha)\,K_{\ell}\big(x,K_{L}(x)\big)+\alpha\,K_{nl}\big(x,K_{L}(x)\big),$
where $\alpha\in[0,1]$ is trainable, $K_{\ell}$ is the linear head, and $K_{nl}$ the nonlinear correction.
The HF loss balances accuracy, linearity preference, and regularization:
$\mathcal{L}_{HF}=\frac{1}{N_{HF}}\sum_{j=1}^{N_{HF}}\!\big(K_{H}(x_{j})-f_{H}(x_{j})\big)^{2}+\lambda_{\alpha}\,\alpha^{\,n}+w\,\|\Phi_{nl}\|_{2}^{2}.$
Here $\lambda_{\alpha}>0$ and $n\in\mathbb{N}$ control the preference for the linear block, while $w>0$ regularizes the nonlinear one.
The intuition is that the LF model captures the dominant structure, while the HF nonlinear head corrects only the residual. This reduces overfitting and stabilizes training when HF data are scarce. MFKAN achieved lower relative $L_{2}$ errors than single-fidelity KANs under the same HF budget [189]. The same fidelity split also benefits PIKANs, where $K_{L}$ serves as a coarse physics surrogate and the HF head is trained only on residual or boundary losses, making fidelity decomposition complementary to spatial domain decomposition (Sec. 7.3).
In a related direction, Jacob et al. [190] [^39] https://github.com/pnnl/spikans propose a separable PIKAN (SPIKAN), expressing the solution as a sum of products of 1D KAN factors—one per coordinate—before combining them into a global field. Bühler et al. [231] propose KAN-SR, a symbolic regression framework inspired by Udrescu et al. [230]. The method checks for separability or symmetry before modeling. If such structure exists, the function is decomposed into smaller subproblems, each fit with a compact KAN, and then recomposed into a closed-form symbolic expression. This divide–and–conquer strategy improves interpretability and robustness in noisy or high-dimensional settings.
@@
7.5 Hybrid, Ensemble, and Data-Integrated Models
A complementary line of work boosts accuracy not by modifying KAN bases, but by combining KAN modules with other neural components—MLPs, attention blocks, or ensemble mechanisms. These hybrids exploit KAN’s structured 1D functional mappings while leveraging the expressive power of deep networks.
  • MLP–KAN hybrids. These architectures use MLPs for broad feature extraction and KANs for structured refinement. Examples include the MLP–KAN mixture-of-experts of He et al. [41] and the HPKM-PINN parallel fusion model of Xu et al. [58].
  • Task-specific KAN replacements. Kundu et al. [257] (KANQAS) [^40] https://github.com/Aqasch/KANQAS_code replace the MLP in a Double Deep Q-Network with a KAN module for quantum architecture search.
  • Structured functional hybrids. KKAN [52] inserts small per-dimension MLPs to generate learned one-dimensional features $\Psi_{p,q}(x_{p})$, sums them across dimensions, and applies a simple outer basis expansion (Chebyshev, Legendre, sinusoidal, or RBF). This yields a pipeline
    $\text{inputs}\rightarrow\text{inner MLPs}\rightarrow\text{dimensionwise sum}\rightarrow\text{basis expansion}\rightarrow\text{output},$
    combining expressive learned features with a stable, low-bias basis. KKAN reports consistent gains over vanilla KANs and MLPs across regression, operator learning, and PDE benchmarks (Fig 19).
Figure 19: KKAN architecture (adapted from [52]). Each input coordinate $x_{p}$ These are summed across dimensions to form $\xi_{m}$, which are expanded via basis functions in the outer block. A final summation yields the output. Bases may include Chebyshev/Legendre polynomials, sine series, or RBFs. This “MLP inside + basis outside” layout mirrors the Kolmogorov–Arnold recipe.
7.6 Sequence and Attention Hybrids
KANs are increasingly combined with sequence encoders and attention mechanisms to capture long-range temporal structure, global context, or multi-scale interactions. Most architectures fall into three categories: (i) attention-enhanced KANs, (ii) sequence-to-field encoder–decoder hybrids, and (iii) generative or data-augmentation models that stabilize training.
  • Attention-enhanced KANs. These models insert KAN blocks into attention layers to improve multi-scale representation. Includes KAN–MHA [56], FlashKAT (Group-Rational KAN in Transformers) [193], and Attention-KAN-PINN for battery SOH forecasting [207].
@@ 7.6 Sequence and Attention Hybrids
  • Generative / data-augmentation hybrids. These integrate KANs with GAN or adaptive-activation mechanisms to improve learning from sparse or noisy data. Includes AAKAN–WGAN [195], which uses a Wasserstein GAN to augment training data.
Across all designs, the auxiliary module supplies memory, context, or synthetic data, while the KAN block provides a compact and interpretable functional mapping at the output stage.
7.7 Discontinuities and Sharp Gradients
KAN variants approach discontinuities, shocks, and steep gradients through three main mechanisms: (i) using bases that naturally represent non-smooth features, (ii) attention or residual reweighting that concentrates learning where gradients are large, and (iii) physics-driven stabilization for shock-dominated PDEs.
  • Non-smooth or localized bases. Designed to capture kinks, singularities, or asymptotic regimes. Includes SincKAN [49], rKAN [128], and KINN for stress singularities [69].
  • Attention and adaptive residuals. Focuses computational effort on regions with steep gradients. Includes KAN–MHA [56], KKAN+ssRBA [52], and AIVT [221].
  • Physics-driven shock stabilization. Embeds stabilization mechanisms from numerical PDEs directly into the loss. Includes cPIKAN+EVM/RBA [37], Buckley–Leverett KANs [61], and EPi–cKAN for elasto-plastic transitions [62].
Overall, these approaches mitigate non-smoothness by introducing localized bases (Sinc, rational), adaptive residual focusing, or physics-aware stabilization that controls oscillations near shocks.
Edge-aware KANs. Lei et al. [60] embed an explicit edge activation inside each KAN layer, enabling sharp jumps while keeping the rest of the representation smooth:
$\psi(x)=w_{t}\,\tanh\!\big(\alpha(x-\beta)\big)+w_{s}\sum_{i}c_{i}\,S_{i}(x),$
where $\alpha,\beta,w_{t},w_{s},c_{i}\in\mathbb{R}$ and $S_{i}(\cdot)$ are B-spline basis functions.
The $\tanh$ term places and sharpens the discontinuity, while the spline component fits the surrounding smooth regions. Incorporated into a PIKAN framework, this mechanism reduces spectral bias and yields more stable representations of steep fronts, a persistent difficulty for both MLPs and classical spline-based KANs.
7.8 Optimization & Adaptive Training
Training KANs—especially in physics-informed or data-scarce settings—relies heavily on optimization choices. Two themes dominate: (i) hybrid optimization schedules and (ii) Bayesian/probabilistic methods.